• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dalton

    Log entry time 18:34:16 on October21,2009

    Entry number 296019

    This entry is a followup to: 295982


    keyword=Laser status

    The laser pointing has stabilised and entered its usual 2 hour pointing cycle. The cavity shows no sign of locking as is. The QPD settings are different from the last time I took note on 30 September - before the last time we had to realign manually (see log entry 292561.)
    I will steer the beam until these positions are duplicated and see if we lock.

    FIGURE 1